16) Saxbury Corporation's relevant range of activity is 3,000 units to 7,000 units. When it produces and sells 5,000 units, its average costs per unit are as follows:

Average Cost per Unit

Direct materials

$5.30

Direct labor

$3.65

Variable manufacturing overhead

$1.50

Fixed manufacturing overhead

$3.90

Fixed selling expense

$0.75

Fixed administrative expense

$0.60

Sales commissions

$0.50

Variable administrative expense

$0.50

Required:

f. If 6,000 units are produced, what is the total amount of fixed manufacturing cost incurred?

g. If 6,000 units are produced, what is the total amount of manufacturing overhead cost incurred? What is this total amount expressed on a per unit basis?

h. If the selling price is $22.90 per unit, what is the contribution margin per unit sold?

i. If 4,000 units are produced, what is the total amount of direct manufacturing cost incurred?

j. If 4,000 units are produced, what is the total amount of indirect manufacturing cost incurred?

k. What incremental manufacturing cost will the company incur if it increases production from 5,000 to 5,001 units?
Answer #1
f
Total amount of fixed manufacturing cost 19500 =5000*3.9
g
Variable manufacturing overhead 9000 =6000*1.5
Fixed manufacturing overhead 19500 =5000*3.9
Total amount of manufacturing overhead cost 28500
Manufacturing overhead per unit 4.75 =28500/6000
h
Selling price 22.90
Less: Variable costs
Direct materials 5.30
Direct labor 3.65
Variable manufacturing overhead 1.50
Sales commissions 0.50
Variable administrative expense 0.50
Total Variable costs 11.45
Contribution margin per unit sold 11.45
i
Direct materials 21200 =4000*5.3
Direct labor 14600 =4000*3.65
Total amount of direct manufacturing cost 35800
j
Variable manufacturing overhead 6000 =4000*1.5
Fixed manufacturing overhead 19500 =5000*3.9
Total amount of indirect manufacturing cost 25500
k
Direct materials 5.30
Direct labor 3.65
Variable manufacturing overhead 1.50
Incremental manufacturing cost 10.45